Tohban Report 2016-11-02
Tohban Reports | |
---|---|
Start Date: | 26 October 2016 |
End Date: | 02 November 2016 |
Tohban: | Pascal Saint-Hilaire |
Tohban email: | pascal@ssl.berkeley.edu |
Next Tohban: | Nicole |
List all reports |
Solar Activity
The solar activity was very low: a few low-B flares. A large coronal hole was present at the beginning of the tohban week, but has since rotated off the disk. Low (6%) probability of >C flare activity in the next day, according to solar monitor.
How many GOES flares occurred?
Flares above B, C, M, X class were 13 0 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 2 0 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 2 0 0 0
There were RHESSI flares/GOES flares 16 / 3 over the time range 18-Oct-16 25-Oct-16
Memory Management
The SSR has been creeping up: now about 30% at end of pass set. Presumably due to the fact that we have slighly less passes a day than before (we only telemeter during RHESSI day now). I've asked ops to get some Waldheim passes to compensate.
Spacecraft Status
D9f livetime is slowly decreasing, now at 80% level. D9r livetime has been plummeting, even with an increase in fast thresholds from 0xA0 to 0XB0. Recommend to increase front by 0x10 (to 0xA0) and rear by 0x20 (to 0xD0).
There is only one BGS pass per day these days (the rest are AGO), so it is a bit more difficult to closely monitor the detectors. After next Friday, there are no planned BGS passes for almost 10 days!
Cold tips have increased ~2.5 K over this week (to ~120K and ~118K), and cold plates by ~3 K (to 141 K for CP2). Shadow length is about 35 mins, and is expected to remain constant over the next couple of weeks.
Data Gaps
Very large data gaps on Nov 1 and Oct 31 -- probably not processed yet.
Some important gaps:
GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 2 GAP START TIME GAP END TIME GAP (SEC) 2016-10-28T02:10:00.000 -- 2016-10-28T02:15:00.000 300.00000 2016-10-28T05:00:00.000 -- 2016-10-28T06:50:00.000 6600.0000
Jimm is looking into that one.
Detector issues
Detectors 3 and 9 are powered on, and are recording events. The D9 rear fast threshold was raised from 0xA0 to 0xB0.
2016-306-21:03:28 /IDPUTABLE9 OFFSET=REARFASTDAC 2016-306-21:03:40 /IDPULOAD VALUE=0xB0
Notes
To spare the battery, data is telemetered only when spacecraft is in sunlight, severely reducing our ability to quickly monitor any effect of e.g. threshold changes, as ground passes with other stations than BGS typically do not give us monitor rates in real time.
Spacecraft Management
Decimation | Normal/Vigorous |
HLAT Decimation | No decimation |
Night time data (fronts) | full-time |
Night time data (rears) | full-time |
Require extra passes? | Yes |
Requirement for moving pointer? | No |
Attenuator operation | Working as expected |
Detector problems? | See notes above. |